# ColorwAI: Generative Colorways of Textiles through GAN and Diffusion Disentanglement

### v1

@@ Abstract.
Abstract. Colorway creation is the task of generating textile samples in alternate color variations maintaining an underlying pattern. The individuation of a suitable color palette for a colorway is a complex creative task, responding to client and market needs, stylistic and cultural specifications, and mood. We introduce a modification of this task, the ”generative colorway” creation, that includes minimal shape modifications, and propose a framework, ”ColorwAI”, to tackle this task using color disentanglement on StyleGAN and Diffusion. We introduce a variation of the InterfaceGAN method for supervised disentanglement, ShapleyVec. We use Shapley values to subselect a few dimensions of the detected latent direction. Moreover, we introduce a general framework to adopt common disentanglement methods on any architecture with a semantic latent space and test it on Diffusion and GANs. We interpret the color representations within the models’ latent space. We find StyleGAN’s W space to be the most aligned with human notions of color. Finally, we suggest that disentanglement can solicit a creative system for colorway creation, and evaluate it through expert questionnaires and creativity theory.
Colorways, Disentanglement, Textiles, Diffusion, GANs
copyright: acmcopyrightjournalyear: 2024doi: XXXXXXX.XXXXXXXccs: Applied computing Arts and humanitiesccs: Applied computing Media artsccs: Computing methodologies Computer vision
1. Introduction
@@
Color profoundly impacts the general sentiment expressed by a textile, influencing its perceptual structure and depth. Its importance is mirrored in the design brief preceding everyday clothing and furniture production. Alongside target specifications, technical requirements, and suggested moods, a crucial part of this brief indicates the desired colorways (Briggs-Goode and Townsend, 2011; Dickinson, 2011), alternate color variations of the same pattern. The market typically demands around 6-8 colorways to meet diverse societal needs.
@@
These colorways can range from slight base-color variations that maintain the original mood to completely different colorations that alter the feel, as in Figure 1. The creation of textile colorways is complex and subjective, influenced by historical and cultural contexts, technical discoveries, trends, seasonal requirements, and individual preferences (Dickinson, 2011).
@@
Figure 1. Example of a colorway in cold and warm tones. Pattern generated using StyleGAN2-ADA and colorway generated using InterfaceGAN.
[…]
Example of a colorway in cold and warm tones. Pattern generated using StyleGAN2-ADA and colorway generated using InterfaceGAN.
@@
With traditional graphics programs and existing AI re-colorization methods, to change the color scheme of the graphic design, a color palette must be predetermined by the designer. With ColorwAI, the framework proposed in the paper, the designer is invited to ”brainstorm with the machine about the desired result”, tackling the ideation moment of the colorway. The designer proposes a color and ColorwAI suggests the pattern in a novel color scheme and, so on, through a repeating dialogue. In this way, the model prompts the designer to consider original possibilities, sometimes adjusting the design’s shapes. The model, because of the knowledge gained from learning historical textiles, adjusts the pattern with slightly different lines and shapes, to be more suitable to the new color scheme identified. It changes the shape of the pattern to the minimum, in a more harmonious way in the model’s lens.
@@
The purpose of the model is to suggest a ”generative colorway”, a colorway that generates both chromatic and slight pattern variation, to stimulate the designer’s creativity. This is intended to accompany the creative process of the textile expert, who is traditionally given minimal specifications from the design brief and must present colorways that are stylistically and conceptually fitting the original design, but novel and original.
@@
We propose to tackle the task of ”generative colorway” creation with our framework ColorwAI, by relying on disentanglement. In deep neural networks, acquired representations are high-dimensional latent spaces encompassing a large variety of features. Disentanglement refers to the extraction of an interpretable feature such as color from the high dimensional latent space. Disentanglement learning is the task of learning a meaningful traversal direction in the latent space corresponding to a semantic variation of the output (Choi et al., 2021; Bengio, 2013). A direction that only influences one semantic factor is considered disentangled from all other factors of variation. In the context of machine learning and computer vision, disentanglement is a well-established method for image editing and interpretation of the hidden representations of generative neural networks.
@@
Because of its nature, weakly supervised disentanglement can determine the representation of color as a dimension within the latent space of generative neural networks (Locatello et al., 2020). The color can be edited by perturbing along the latent direction, thus creating a ”generative colorway” around that base color. Because generative neural networks such as VAEs, GANs, and Diffusion feature at least one coherently and semantically arranged latent space (Karras et al., 2019; Kwon et al., 2022; Higgins et al., 2016), we assume that a disentangled color direction provides visually plausible color modifications, responding to the need for a suitable palette for the textile. Whenever a perturbation from a starting pattern along the latent direction of a color incurs into a lesser defined, warped, area of the latent space, the modifications are more substantial, incurring structural modifications. This may occur for a textile pattern when a certain chromaticity, e.g. blue, is lacking in the training data. In such cases, disentanglement creates slight pattern modifications, adjusting the original textile pattern into a similar one for which a blue example does exist in the training data.
@@
An additional advantage of disentanglement is that it allows understanding how colors are represented within these models in the context of textiles. Given the wealth of cultural, societal, and technical factors behind the use of color in textiles, as expanded in Section 2.1, we wish to complement the practical framework with insights into how the methods that generate these colorways represent colors.
@@
In this paper, we present ColorwAI, a framework for ”generative colorway” creation that can be flexibly applied to any generative architecture with a semantic latent space. We identify four stages in the disentanglement pipeline: the coupling of latent codes with image annotations, the disentanglement of the latent space, the modifications of the latent code, and the inference with modified latent code. We show disentanglement results for StyleGAN and DDPM (Karras et al., 2019; Ho et al., 2020), maintaining distinct coupling and inference steps, but common disentanglement methods. This approach to Diffusion is a generalization of BoundaryDiffusion (Zhu et al., 2024). We test InterfaceGAN and StyleSpace disentanglement methods and introduce a modification of InterfaceGAN, that we call ShapleyVec. The latter uses a subset of latent space neurons for disentanglement (Wu et al., 2021; Shen et al., 2020). We demonstrate the superiority of the latter method in terms of quantitative performance and interpretability.
@@
  1. (1)
    We introduce a new task motivated by practical utility, the ”generative colorway” creation, that extends existing work on color re-colorization with creative controls;
  2. (2)
@@
2. Context
2.0.1. Textiles and Colors
Colors and textures create complex interactions influencing human perception and artistic evolution. Textiles, originally referring to woven (from Latin textilis) fabrics, have accompanied humans since the beginning of civilization, becoming a distinctive factor of humanity. In the West European-Mediterranean sphere, its history interlaced with that of color already 5000 years ago in Egypt (Whiston, 1948). The colors came from the fibers and dyes. Their pigments were extracted from elements of nature, as in the ancient tradition of the madder root, creating shades of red and orange, and the expensive carmine-yielding cochineal insects.
Pigment availability influenced cultural identities. Christian and Copts [^1] The Copts are an indigenous Christian ethnoreligious community in Egypt. colors are distinctively subdued in line with religious and social reasons, as opposed to the bright pigments of Assyrians and Babylonians, mirroring their active and strenuous civilizations but also the availability of the vibrant Tyrian purple (Whiston, 1948). In more recent centuries, color developments instantiated trends that characterized their time, such as mauveine and analine purple in the 1850s and turquoise a century later (Dickinson, 2011).
@@ 2.0.1. Textiles and Colors
The use of color in textiles is influenced by its perception. Studies initiated from Chevreul’s ’Laws of Contrast of Colour’ have shown that the co-location of colors into several schemes creates a harmonious product (Burchett, 2002), becoming the de-facto standard in color design. Bauhaus artists such as Johannes Itten and Josef Albers, influenced by Goethe’s ’Theory of Colors’, explored color relationships affecting human perception. Josef Albers (1963) investigated color interaction. He demonstrated how perception of two identical colors on different backgrounds yields perceptually distinct colors (Albers, 2013). Rules and costumes were introduced after the Great Exhibition in London of 1851, including the ’Grammar of Ornaments’ by Owen Jones (Jones, 2016). Jones mentions, in the 37 principles for textile design, the importance of using a pondered distribution of hues on the surface.
@@
Using digital synthetic textures, (Wang et al., 2020) demonstrates the influence of shape on the perception of the color composition in patterns, showing how more elongated shapes lead to less prominent chromatic impressions. The finding was already experimentally clear to Albers and Jones. Our perception of color similarity and arrangement is merely functional. Particularly, color patterns are very modestly understood by our vision compared to texture, contrast, and color (Mojsilovic et al., 2000).
2.0.2. Texture Generation and Re-colorization
@@
Relevant research in Computer Vision has focused on the generation of textures (Portilla and Simoncelli, 2000; Karagoz et al., 2023; Dai et al., 2014; Lin et al., 2023). By texture synthesis, we mean the generation of images with structural components, spatially homogeneous and repeating, often fabrics or materials samples. Early works adopt parametric statistical models such as Portilla et al. (Portilla and Simoncelli, 2000), using complex wavelet coefficients. More recently, papers use deep learning-based generative models such as AE, GANs, and Diffusion. Bergmann et al. (Bergmann et al., 2018) introduce a structural similarity loss to detect textural defects using Autoencoders. Lin et al. (Lin et al., 2023) exploits StyleGAN to generate realistic visual textures, obtaining high-quality samples and inversion (Karras et al., 2019). Lastly, Karagoz et al. (Karagoz et al., 2023) generate textile patterns using Stable Diffusion conditioned with a number of keywords from textile metadata (Rombach et al., 2022).
@@
Given the success of the last two models and the proven semantic arrangement of their latent space (Karras et al., 2019; Kwon et al., 2022), we generate textile samples and colorways using StyleGAN and Diffusion.
@@
The generation of colorways is closely linked with the task of image re-colorization. In this context, Xu et al. (Xu et al., 2019) developed a re-colorization algorithm on fabrics based on parts segmentation and variation based on given color themes. Cho et al. (Cho et al., 2017) introduced the first deep neural network for content-aware image re-colorization. Yang et al. (Yang et al., 2021), among other factors, disentangle color mood within scene depictions. The color control is only partially addressed in (Yang et al., 2021). In the context of conditioning methods, Li et al. (Li et al., 2020) use text conditioning to vary the color of semantic image segments using natural language.
@@
2.0.3. Disentanglement
Recent breakthroughs in deep neural networks for image generation have significantly impacted the artistic and design industries  (Kingma and Welling, 2013; Goodfellow et al., 2020; Rombach et al., 2022), enhancing image editing and manipulation through disentanglement and text-conditioning techniques  (Gatys et al., 2016; Karras et al., 2019; Patashnik et al., 2021; Wang et al., 2023). This study examines the creative and interpretative aspects of these advancements.
Methodologies for determining and interpreting the latent factors of variation have been developed starting from variations of VAE and GAN models in supervised and unsupervised settings (Karras et al., 2019; Higgins et al., 2016). Harkonen et al. (Härkönen et al., 2020), for instance, discover interpretable image controls using Principal Component Analysis (PCA) on the latent space of the model. Using supervision, Shen et al. (Shen et al., 2020) and Yang et al. (Yang et al., 2021) train a machine learning regression to find the separation boundary between different semantics in the latent space. Methods such as Wu et al. (Wu et al., 2021) focus, instead, on the identification of specific dimensions and channels that disentangle the selected feature.
Recently, attention has been dedicated to disentanglement in Diffusion models. This task is more complex than the disentanglement of earlier methods as Diffusion models do not present a unique deterministic latent space. Some works focus on the UNet bottleneck $h-space$ or latent representation $\epsilon-space$ using deterministic variations of the scheduler and inversion (Kwon et al., 2022; Zhu et al., 2024), while others disentangle the text conditioning from CLIP (Wu et al., 2023).
3. ColorwAI
This section introduces the ”generative colorways” framework, detailing the StyleGAN and DDPM models, their latent space properties, and disentanglement techniques. We describe methods for discovering latent traversal directions, including InterfaceGAN, StyleSpace, and our modified ShapleyVec (Fig.  2).
Figure 2. Schematization of the ColorwAI framework. We show the disentanglement steps for the two architectures, their latent spaces, and the inference procedure. Steps 1 and 4 are different for GANs and Diffusion, while 2 and 3 are common.
[…]
Schematization of the ColorwAI framework. We show the disentanglement steps for the two architectures, their latent spaces, and the inference procedure. Steps 1 and 4 are different for GANs and Diffusion, while 2 and 3 are common.
3.1. Latent Spaces
3.1.1. StyleGAN
StyleGAN2-ADA is a style-based GAN generative neural network belonging to the family of StyleGANs (Karras et al., 2019). The ADA version is more suitable for training with smaller data than the original and versions 2 and 3 of StyleGAN (Karras et al., 2021).
@@ 3.1.1. StyleGAN
3.1.2. Diffusion Denoising Probabilistic Models
In this section we outline the core principles of Diffusion models (Ho et al., 2020), leaving in the appendix more details.
Diffusion models (Ho et al., 2020) generate data by learning how to reverse a forward diffusion process that gradually adds noise to the data. Starting from pure noise, the reverse process progressively refines the image to the desired outcome.
@@ 3.1.2. Diffusion Denoising Probabilistic Models
(1) $q(\mathbf{x}t|\mathbf{x}{t-1})=\mathcal{N}(\mathbf{x}t;\sqrt{1-\beta_{t}}\mathbf{x}{t-1},\beta_{t}\mathbf{I})$
@@
(2) $p_{\theta}(\mathbf{x}{t-1}|\mathbf{x}t)=\mathcal{N}(\mathbf{x}{t-1};\mu\theta(\mathbf{x}t,t),\Sigma\theta(\mathbf{x}_{t},t))$
DiffusionCLIP integrates diffusion models with CLIP to guide the generation process using text embeddings, by conditioning the reverse process (Kim et al., 2022; Radford et al., 2021).
Asyrp propose to modify the generative reverse process to use the deepest feature maps, the h-space, of a pre-trained diffusion model as semantic latent space for image manipulation, using a modified deterministic DDIM denoising schedule.
Boundary Diffusion segments the latent space into distinct regions for precise attribute control without additional parameters (Zhu et al., 2023). This method modifies the reverse process to include boundaries during the mixing step, which is the denoising step where the final image starts to shape up. This method finds the disentangled direction of a feature by determining its semantic boundary using SVM. The latent direction is the orthonormal direction exiting this boundary, similar to the InterfaceGAN method used for GANs (Shen et al., 2020). This indicates that other GAN disentanglement methods, such as StyleSpace and ShapleyVec, can also be applied within the Boundary Diffusion framework.
3.2. Disentanglement Methods

### v2

@@ Abstract
Colorway creation is the task of generating textile samples in alternate color variations maintaining an underlying pattern. Selecting a colorway is a complex creative task, responding to client and market needs, technical and cultural specifications, and personal artist style. We introduce a framework, "ColorwAI", to tackle the generative task using color disentanglement on StyleGAN and Diffusion while maintaining minimal shape alteration. We present a variation of the InterfaceGAN method for semi-supervised disentanglement, ShapleyVec, which uses Shapley values to subselect salient dimensions from the detected latent direction. Moreover, we present a framework to employ common disentanglement methods on any architecture with a semantic latent space, and test it on DDM and StyleGAN2-ADA. Our results show that StyleGAN’s W space is the most aligned with human notions of color in terms of vector similarities and generated colorways. Finally, we suggest that disentanglement can solicit a creative system for colorway creation, and evaluate it through expert questionnaires and within the lens of creativity theory.
Keywords:
Colorways Disentanglement Textiles Diffusion GANs
1 Introduction
@@
Color profoundly impacts the general sentiment expressed by a textile, influencing its perceptual structure and depth. Its importance is mirrored in the design brief preceding everyday clothing and furniture production. Alongside target specifications, technical requirements, and suggested moods, a crucial part of this brief indicates the desired colorways [7, 12], alternate color variations of the same pattern. The market typically demands around 6-8 colorways to meet diverse societal needs [7].
@@
These colorways can range from slight base-color variations that maintain the original tone to completely different colorations that alter the feel, as in Figure 1. The creation of textile colorways is complex and subjective, influenced by historical and cultural contexts, technical discoveries, trends, seasonal requirements, and individual preferences [12].
@@
Figure 1: Example of a colorway in cold and warm tones. Pattern generated using StyleGAN2-ADA and colorway generated using InterfaceGAN.
@@
With traditional graphics programs and existing AI re-colorization methods, a color palette must be predetermined by the designer to change the color scheme of the graphic design. With ColorwAI, the framework proposed in the paper, designers and artists are assisted in the creative process, in the ideation of the colorway. The designer proposes a color and ColorwAI suggests the pattern in a novel color scheme and, so on, iteratively. In this way, the model prompts the designer to consider original possibilities, sometimes adjusting the design’s shapes. The model, differently from re-colorization methods, adjusts the pattern slightly, to be more suitable to the new color scheme identified. It changes the shape of the pattern to the minimum, in a more harmonious way according to the knowledge gained from training on historical textiles.
@@
The purpose of the model is to suggest a "generative colorway", a colorway that generates both chromatic and slight pattern variation, to stimulate the designer’s creativity. This is intended to accompany the creative process of the textile expert, who is traditionally given minimal specifications from the design brief and must present colorways that are stylistically and conceptually fitting the original design, but novel and original.
@@
ColorwAI tackles this creative process by relying on disentanglement. In deep neural networks, acquired representations are high-dimensional latent spaces encompassing a large variety of features. Disentanglement refers to the extraction of an interpretable feature such as color from the high dimensional latent space. Disentanglement learning is the task of learning a meaningful traversal direction in the latent space corresponding to one semantic variation of the output [10, 2]. The latter is a well-established method for image editing and interpretation of the hidden representations of generative neural networks.
@@
Because of its nature, weakly supervised disentanglement can determine the representation of color as a dimension within the latent space of generative neural networks [30]. The color can be edited by perturbing along the latent direction, thus creating a "generative colorway" around that base color. Because generative neural networks such as VAEs, GANs, and Diffusion feature at least one coherently and semantically arranged latent space [22, 26, 16], we assume that a disentangled color direction provides visually plausible color modifications, responding to the need for a suitable palette for the textile. Whenever a perturbation from a starting pattern along the latent direction of a color incurs into a lesser defined, warped, area of the latent space, the modifications are more substantial, incurring structural modifications. This may occur for a textile pattern when a certain chromaticity, e.g. blue, is lacking in the training data. In such cases, disentanglement creates slight pattern modifications, adjusting the original textile pattern into a similar one for which a blue example does exist in the training data.
@@
An additional advantage of disentanglement is that it allows understanding how colors are represented within these models in the context of textiles. Given the wealth of cultural, societal, and technical factors behind the use of color in textiles, we wish to complement the practical framework with insights into how our methods represent colors.
@@
In this paper, we present ColorwAI, a framework for "generative colorway" creation that can be flexibly applied to any generative architecture with a semantic latent space. We identify four stages in the disentanglement pipeline: the coupling of latent codes with image annotations, the disentanglement of the latent space, the modifications of the latent code, and the inference with modified latent code. We show disentanglement results for StyleGAN and DDM [22, 17], maintaining distinct coupling and inference steps, but common disentanglement methods. This approach to Diffusion is a generalization of BoundaryDiffusion [52]. We test InterfaceGAN and StyleSpace disentanglement methods and introduce a modification of InterfaceGAN, that we call ShapleyVec. The latter uses a subset of latent space neurons for disentanglement [48, 41]. We demonstrate the superiority of the latter method in terms of quantitative performance and interpretability.
@@
  1. 1.
    We introduce a new task motivated by practical utility, the "generative colorway" creation, that extends existing work on re-colorization with creative controls;
  2. 2.
@@
2 Context
2.0.1 Textiles and Colors
Colors and textures create complex interactions influencing human perception and artistic evolution. Textiles, originally referring to woven (from Latin textilis) fabrics, have accompanied humans since the beginning of civilization, becoming a distinctive factor of humanity. In the West European-Mediterranean sphere, its history interlaced with that of color already 5000 years ago in Egypt [46]. The colors came from the fibers and dyes. Their pigments were extracted from elements of nature, as in the ancient tradition of the madder root, creating shades of red and orange, and the expensive carmine-yielding cochineal insects.
Pigment availability influenced cultural identities. Christian and Copts [^1] The Copts are an indigenous Christian ethnoreligious community in Egypt. colors are distinctively subdued in line with religious and social reasons, as opposed to the bright pigments of Assyrians and Babylonians, mirroring their active and strenuous civilizations but also the availability of the vibrant Tyrian purple [46]. In more recent centuries, color developments instantiated trends that characterized their time, such as mauveine and analine purple in the 1850s and turquoise a century later [12].
@@ 2.0.1 Textiles and Colors
The use of color in textiles is influenced by its perception. Studies pioneered by Chevreul’s Laws of Contrast of Colour have shown that the co-location of colors into several schemes creates a harmonious product [8], becoming the de-facto standard in color design. Bauhaus artists such as Johannes Itten and Josef Albers, influenced by Goethe’s Theory of Colors, explored color relationships affecting human perception. Josef Albers (1963) investigated color interaction. He demonstrated how perception of two identical colors on different backgrounds yields perceptually distinct colors [1]. Rules and costumes were introduced after the Great Exhibition in London of 1851, including the ’Grammar of Ornaments’ by Owen Jones [18]. Jones mentions, in the 37 principles for textile design, the importance of using a pondered distribution of hues on the surface.
@@
Using digital synthetic textures, [44] demonstrates the influence of shape on the perception of the color composition in patterns, showing how more elongated shapes lead to less prominent chromatic impressions. The finding was already experimentally clear to Albers and Jones. Our perception of color similarity and arrangement is merely functional. Particularly, color patterns are very modestly understood by our vision compared to texture, contrast, and color [32].
2.0.2 Texture Generation and Re-colorization
@@
Relevant research in Computer Vision has tackled the generation of textures [35, 20, 11, 28]. By texture synthesis, we mean the generation of images with structural components, spatially homogeneous and repeating, often fabrics or materials samples. Early works adopt parametric statistical models such as Portilla et al. [35], using complex wavelet coefficients. More recently, papers have used deep learning-based generative models such as AE, GANs, and Diffusion. Bergmann et al. [3] introduce a structural similarity loss to detect textural defects using Autoencoders. Lin et al. [28] exploits StyleGAN to generate realistic visual textures, obtaining high-quality samples and inversion [22]. Lastly, Karagoz et al. [20] generate textile patterns using Stable Diffusion conditioned with a number of keywords from textile metadata [39].
@@
Given the success of the last two models and the proven semantic arrangement of their latent space [22, 26], we generate textile samples and colorways using StyleGAN and Diffusion.
@@
The generation of colorways is closely linked with the task of image re-colorization. In this context, Xu et al. [49] developed a re-colorization algorithm on fabrics based on parts segmentation and variation based on given color themes. Cho et al. [9] introduced the first deep neural network for content-aware image re-colorization. Yang et al. [50], among other factors, disentangle color mood within scene depictions. The color control is only partially addressed in [50]. In the context of conditioning methods, Li et al. [27] use text conditioning to vary the color of semantic image segments using natural language.
@@
2.0.3 Disentanglement
Recent breakthroughs in deep neural networks for image generation have significantly impacted the artistic and design industries  [24, 14, 39], enhancing image editing and manipulation through disentanglement and text-conditioning techniques  [13, 22, 34, 45]. This study examines the creative and interpretative aspects of these advancements.
Methodologies for determining and interpreting the latent factors of variation have been developed starting from variations of VAE and GAN models in supervised and unsupervised settings [22, 16]. Härkönen et al. [15], for instance, discover interpretable image controls using Principal Component Analysis (PCA) on the latent space of the model. Using supervision, Shen et al. [41] and Yang et al. [50] train a machine learning regression to find the separation boundary between different semantics in the latent space. Methods such as Wu et al. [48] focus, instead, on the identification of specific dimensions and channels that disentangle the selected feature.
Recently, attention has been dedicated to disentanglement in Diffusion models. This task is more complex than the disentanglement of earlier methods as Diffusion models do not present a unique deterministic latent space. Some works focus on the UNet bottleneck $h-space$ or latent representation $\epsilon-space$ using deterministic variations of the scheduler and inversion [40, 52, 42], while others disentangle the text conditioning [47].
3 ColorwAI
This section introduces the "generative colorways" framework, detailing the StyleGAN and DDM models, their latent space properties, and disentanglement techniques. We describe methods for discovering latent traversal directions, including InterfaceGAN, StyleSpace, and our modified ShapleyVec (Fig. 2).
Figure 2: Schematization of the ColorwAI framework. We show the disentanglement steps for the two architectures, their latent spaces, and the inference procedure. Steps 1 and 4 are different for GANs and Diffusion, while 2 and 3 are common.
3.1 Latent Spaces
3.1.1 StyleGAN
StyleGAN2-ADA is a style-based GAN generative neural network belonging to the family of StyleGANs [22]. The ADA version is more suitable for training with smaller data than the original and versions 2 and 3 of StyleGAN [21].
@@ 3.1.1 StyleGAN
3.1.2 Diffusion Denoising Models
In this section we outline the core principles of Diffusion models [17], leaving in the appendix more details.
Diffusion models [17] generate data by learning how to reverse a forward diffusion process that gradually adds noise to the data. Starting from pure noise, the reverse process progressively refines the image to the desired outcome.
@@ 3.1.2 Diffusion Denoising Models
$q(\mathbf{x}t|\mathbf{x}{t-1})=\mathcal{N}(\mathbf{x}t;\sqrt{1-\beta_{t}}\mathbf{x}{t-1},\beta_{t}\mathbf{I})$ (1)
@@
$p_{\theta}(\mathbf{x}{t-1}|\mathbf{x}t)=\mathcal{N}(\mathbf{x}{t-1};\mu\theta(\mathbf{x}t,t),\Sigma\theta(\mathbf{x}_{t},t))$ (2)
DiffusionCLIP integrates diffusion models with CLIP to guide the generation process using text embeddings, by conditioning the reverse process [23, 36].
Asyrp propose to modify the generative reverse process to use the deepest feature maps [26], the h-space, of a pre-trained diffusion model as semantic latent space for image manipulation, using a modified deterministic DDIM denoising schedule [42].
Boundary Diffusion segments the latent space into distinct regions for precise attribute control without additional parameters [51]. This method modifies the reverse process to include boundaries during the mixing step, the denoising step where the final image starts to shape up. This method finds the disentangled direction of a feature by determining its semantic boundary using SVM. The latent direction is the orthonormal direction exiting this boundary, similar to the InterfaceGAN method used for GANs [41]. This indicates that other GAN disentanglement methods, such as StyleSpace and ShapleyVec, can also be applied within the Boundary Diffusion framework.
3.2 Disentanglement Methods